A chemistry graduate student is given 125. mL of a 1.70 M chlorous acid (HNO2) solution. Chlorous acid is a weak acid with Ka= 1.1 x 10^-2. What mass of NaClO2 should the student dissolve in the HClO2 solution to turn it into a buffer with pH = 2.48? You may assume that the volume of the solution doesn't change when the NaClO2 is dissolved in it. Be sure your answer has a unit symbol, and round it to 2 significant digits.
Answer #1

pH of buffer is calculated using Hendersen equation as

pH = pKa + log [conjugate base] /[acid]

given [acid] = 1.70M

and ka of acid = 1.1 x10-2 , thus pKa of acid = 1.96

we need to calculate [conjugate base]

thus substituting the known values

2.48 = 1.96 + log [conjugate base]/ 1.7

ccalclulating we get molarity of NaCLO2 = 5.6287 M

Thus 1000mL of buffer should contain mass of NaClO2 = molarity x molar mass

= 5.6287mol/L x 90.44 g/mol

= 509.06g/L

But our buffer is made of 125mL only.

Thus the mass of NaClO2 to be added to 125mL of buffer = 509.06g/L x 0.125L

= 63.63 g

= 64 g ( rounding to two significant figures)